OpenStudy (anonymous):

What are the coordinates of the vertex of the function below? Write your answer in the form (x,y). y-9=-6(x-1)^2

OpenStudy (anonymous):

The coordinates of the vertex, by vertex form y = a(x - h)² + k, is... y = -6(x - 1)² + 9 Vertex: (1,9) I hope this helps!

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw